Both min and max values?
nge 7.2.13 Find the minimum and maximum values of z 3x+8y, if possible, for the following set of constraints. 4x + 3y 2 12 x+3y 26 x20, y20 rkec Sco ts: Select the correct choice below and, if necessary, fill n the answer box to complete your choice. O A. The minimum value is「, (Round to the nearest tenth as needed.) O B. There is no minimum value.
Answer #1

here constraints have \geq sign so function has only minimum value for maximum value solution is unbound

find minimum value

Min\:\:\: Z = 3 x_1 + 8 x_2 + 0 S_1 + 0 S_2 + M A_1 + M A_2

subject to

4 x_1 + 3 x_2 - S_1 + A_1 = 12

x_1 + 3 x_2 - S_2 + A_2 = 6

Iteration-1

Cj

3

8

0

0

M

M

B

CB

XB

x1

x2

S1

S2

A1

A2

MinRatio
XB/x2

A1

M

12

4

3

-1

0

1

0

12/3=4

A2

M

6

1

(3)

0

-1

0

1

6/3=2

Z=18M

Zj

5M

6M

-M

-M

M

M

Zj-Cj

5M-3

6M-8↑

-M

-M

0

0

Positive maximum Zj-Cj is 6M-8 and its column index is 2

Minimum ratio is 2 and its row index is 2.

The pivot element is 3.

Entering =x2, Departing =A2

R_2 \leftarrow R_2 *\frac{1}{3}

R_1 \leftarrow R_1-3R_2

Iteration-2

Cj

3

8

0

0

M

B

CB

XB

x1

x2

S1

S2

A1

MinRatio
X
B/x1

A1

M

6

(3)

0

-1

1

1

6/3=2

x2

8

2

1/3

1

0

-1/3

0

2/(1/3)=6

Z=6M+16

Zj

3M+8/3

8

-M

M-8/3

M

Zj-Cj

3M-1/3↑

0

-M

M-8/3

0

Positive maximum Zj-Cj is 3M-1/3 and its column index is 1

Minimum ratio is 2 and its row index is 1

The pivot element is 3.

Entering =x1, Departing =A1

R_1 \leftarrow R_1 *\frac{1}{3}

R_2 \leftarrow R_2 -R_1 *\frac{1}{3}

Iteration-3

Cj

3

8

0

0

B

CB

XB

x1

x2

S1

S2

MinRatio

x1

3

2

1

0

-1/3

1/3

x2

8

4/3

0

1

1/9

-4/9

Z=50/3

Zj

3

8

-1/9

-23/9

Zj-Cj

0

0

-1/9

-23/9

all Zj-Cj\leq 0

optimal solution is arrived

{\color{Red} x_1=2,\:\:\:x_2=\frac{4}{3} }

{\color{Red}Min \:\:\:Z=\frac{50}{3} }